# Optical absorption in tilted geometries as an indirect measure of longitudinal plasma waves in layered cuprates

### v1

@@ Abstract
Electromagnetic waves propagating in a layered superconductor with arbitrary momentum with respect to the main crystallographic directions display an unavoidable mixing between longitudinal and transverse degrees of freedom. Here we show that this basic physical mechanism explains the emergence of a well-defined absorption peak in the in-plane optical conductivity for light propagating at small tilting angles with respect to the stacking direction in layered cuprates. More specifically, we show that this peak, often interpreted as a spurious leakage of the $c$-axis Josephson plasmon, is instead a signature of the true longitudinal plasma mode occurring at larger momenta. By combining a classical approach based on Maxwell’s equations with a full quantum derivation of the plasma modes based on the modelling of the superconducting phase degrees of freedom, we provide an analytical expression for the absorption peak as a function of the tilting angle and light polarization. We suggest that an all-optical measurement in tilted geometry can be used as an alternative way to access plasma-wave dispersion, usually measured by means of large-momenta scattering techniques like RIXS or EELS.
@@ I Introduction
In superconductors the breaking of the continuous gauge symmetry below the superconducting (SC) critical temperature is accompanied by the emergence of two collective modes, associated with the amplitude (Higgs) or phase (Goldstone) fluctuations of the complex SC order parameter, whose absolute value at equilibrium defines the spectral gap for single-particle excitations [1]. While the former is a massive excitation, the latter is massless at long wavelength, reflecting the infinity of possible ground states connected by a global change of the order-parameter phase. Nonetheless, the coupling of the SC phase to the density is directly affected by long-range Coulomb interactions between charged electrons. This effect moves the phase mode to the plasma energy scale [2], that is usually much larger than the spectral gap. As a consequence, optical signatures at the plasma energy scale, i.e. at the zero of the dielectric function, are usually unaffected by the SC transition. A rather different phenomenology is instead observed in anisotropic layered superconductors, i.e. systems where the pairing mainly occurs within planes stacked along the $c$ direction, and SC order is established below $T_{c}$ thanks to a weak Josephson-like inter-plane interaction. The hallmark of this category is represented by high-temperature cuprates [3], where the marked anisotropy has been experimentally proven by different optical probes, starting from linear optics, which measures two well-separated energy scales for the plasma modes at long-wavelength for electric fields propagating in the CuO2 planes or perpendicular to them. In these systems the incoherent quasiparticle hopping along the stacking direction makes the $c$-axis response badly metallic: in contrast, below $T_{c}$ the opening of a sizeable spectral gap along with the weak inter-layer pair hopping leave a rather sharp SC plasma edge at a frequency $\omega_{c}$ of few THz in the optical reflectivity, that clearly testifies the emergence of a well-defined SC Josephson plasmon. Even though this feature has been experimentally observed already in the late ’90s [4, 5, 6, 7, 8, 9], a renewed interest in the physics of Josephson plasmons emerged more recently thanks to non-linear THz spectroscopy from one side [10, 11, 12, 13, 14, 15], and to nano-plasmonic from the other side [16, 17]. In both cases it becomes theoretically relevant understanding the momentum dependence of the plasmon dispersion at generic momentum, i.e. not along the main crystallographic axes. In this configuration one immediately realizes that the anisotropy leads to a non-trivial response of the system, due to the fact that the current induced by the external electric field is no more parallel to the field itself. As it has been extensively discussed in details in Ref.s [18, 19, 20], this mechanism leads to a mixing of the longitudinal and transverse response inside the material, making the distinction between plasmons and polaritons blurred at momenta smaller than a scale $\bar{k}\sim\sqrt{\omega_{ab}^{2}-\omega_{c}^{2}}/c$ set by the anisotropy between in-plane $\omega_{ab}$ and out-of-plane $\omega_{c}$ plasma frequencies. Since usually $\omega_{ab}\gg\omega_{c}$, the effect is relevant for non-linear Josephson plasmonics in the THz regime [21, 22, 23, 24], but does not affect e.g. the measurements of plasmons in RIXS [25, 26, 27, 28, 29] or EELS [30, 31, 32], that usually measure momenta in a fraction of the Brillouin-zone. In the present manuscript we investigate an additional consequence of the above-mentioned mixing, showing how even linear optics can be used to disentangle the longitudinal-transverse mixing in a reflection or transmission geometry which highlights the emergence inside the material of a longitudinal response induced by an external transverse electromagnetic wave. The effect manifests as an absorption peak at a scale nearby $\omega_{c}$ for an electromagnetic wave travelling at small angle with respect to the $c$ direction. This feature has been measured in the past in different samples of electron-doped cuprates [33, 34, 35, 36] below $T_{c}$, and it has been often interpreted as a leakage of the $c$-axis plasmon into the in-plane response [37]. Even more interestingly, the peak position has been shown to change by varying the wave polarization in the plane of incidence, challenging considerably the interpretation of the results. Here we provide a full theoretical description of the microscopic mechanism behind the anomalous absorption peak, and we show that it is a direct consequence of the plasmon-polariton mixing in an anisotropic layered superconductor. We argue that this effect can be used to indirectly probe the plasmon dispersion that usually appears in RIXS and EELS experiments at much larger momenta and, by changing the light polarization, to extract the in-plane and out-of-plane plasma frequencies. Our findings are benchmarked against existing experimental data for cuprates. On a more general ground, our results offer a novel perspective on the possibility to access collective polariton modes in complex materials by properly engineering optical measurements.
@@ II Anisotropic linear response of layered system
$\sigma_{t}(\omega,\eta)=\frac{-i\omega\varepsilon_{\infty}/4\pi\big{(}\sigma_{ab}\cos^{2}{\eta}+\sigma_{c}\sin^{2}\eta\big{)}+\sigma_{ab}\sigma_{c}}{-i\omega\varepsilon_{\infty}/4\pi+\sigma_{ab}\sin^{2}\eta+\sigma_{c}\cos^{2}\eta}.$ (4)
For an electric field polarized along $t$ ($\text{E}_{b}=0$), Eq. (4) immediately gives the conductivity we were looking for, $\sigma_{t}=\text{J}_{t}/\text{E}_{t}$. This expression was first derived in Ref. [34], and its real part displays a peak with central frequency that moves with $\eta$. To show it explicitly, we replace $\sigma_{ab}=-\frac{\varepsilon_{\infty}\omega_{ab}^{2}}{4\pi i\omega}$ and $\sigma_{c}=-\frac{\varepsilon_{\infty}\omega_{c}^{2}}{4\pi i\omega}$, where $\omega_{ab}$ and $\omega_{c}$ are the in-plane and out-of-plane plasma frequencies respectively: one then immediately sees that the real part of $\sigma_{t}(\omega,\eta)$ peaks at frequency $\omega_{l}(\eta)$, that reads
@@
$\varepsilon_{ab}(\omega)=\varepsilon_{\infty}\left(1-\frac{\omega_{ab}^{2}}{\omega^{2}}\right),$ (6)
@@
$\varepsilon_{c}(\omega)=\varepsilon_{\infty}\left(1-\frac{\omega_{c}^{2}}{\omega^{2}}\right).$ (7)
@@
$\varepsilon_{t}(\omega,\eta)=\varepsilon_{\infty}\frac{\big{(}\omega^{2}-\omega_{ab}^{2}\big{)}\big{(}\omega^{2}-\omega_{c}^{2}\big{)}}{\omega^{2}\big{(}\omega^{2}-\omega^{2}_{l}(\eta)\big{)}}.$ (8)
@@ III.1 Effective action description of plasma modes
To gain a better physical insight into the results of the previous section, we will take advantage of the description of the plasma modes in the SC state obtained via the phase degrees of freedom. Indeed, as it has been recently discussed in Ref.s [18, 19], this is a rather powerful and elegant approach to describe the interplay between longitudinal and transverse plasma waves in a layered superconductor, that leads to generalized plasma modes with mixed character at low momenta. We summarize here the main ingredients of the derivation, referring the reader to Ref.s [18, 19] and references therein for a detailed derivation of the layered phase-only model.
@@
These gauge-invariant fields provide a full description of the plasma modes once the phase fluctuations are integrated out [18, 19]. In the basis $\boldsymbol{\psi}_{abc}=\begin{pmatrix}\psi_{a}&\psi_{b}&\psi_{c}\end{pmatrix}^{T}$, and by taking the limit for infinite compressibility that is appropriate for cuprates,
@@
where the plasma frequencies are linked to the in-plane and out-of-plane superfluid stiffness, $\omega_{ab}^{2}=4\pi e^{2}D_{ab}/\varepsilon_{\infty}$ and $\omega_{c}^{2}=4\pi e^{2}D_{c}/\varepsilon_{\infty}$ respectively. Indeed, these go back to Eq.s (6) and (7) once the analytic continuation $i\Omega_{m}\to\omega+i0^{+}$ is performed. Notice that the dielectric tensor is diagonal in the basis $\boldsymbol{\psi}_{abc}$, as $(a,b,c)$ is the reference frame of the principal axes of the crystal. By their definition in Eq. (11), the gauge invariant fields are formally proportional to currents, and we can then apply within the effective-action framework the same procedure used above for the classical approach, i.e. a change of the reference frame to describe a transverse dielectric tensor. We thus perform a rotation around the $b$-axis that combines the $\psi_{a}$ and $\psi_{c}$ components into transverse $\psi_{t}$ and longitudinal $\psi_{l}$ components with respect to the momentum k. The matrix that performs the change of basis $\boldsymbol{\psi}_{abc}\to\boldsymbol{\psi}_{tbl}=\begin{pmatrix}\psi_{t}&\psi_{b}&\psi_{l}\end{pmatrix}^{T}$ reads
@@
where we have performed the analytic continuation $i\Omega_{m}\to\omega+i0^{+}$ and used $k_{c}=|\textbf{k}|\cos\eta$ and $k_{a}=|\textbf{k}|\sin\eta$. Using the definitions of the dielectric functions in Eq.s (13) and (14), the solution of Eq. (19) is exactly the frequency
@@ III.2 Interpretation of the conductivity peak of plasmons
is a dielectric function that describes the transverse linear response of the superconductor along the $t$-axis. Indeed, by making use of the relation $\varepsilon_{\alpha}=\varepsilon_{\infty}+4\pi i\sigma_{\alpha}/\omega$ between the optical conductivity and the dielectric function along the direction $\alpha$ [49] one recovers $\sigma_{t}$ as in Eq. (4). Remarkably, the denominator of $\varepsilon_{t}$ can be brought back to the left-hand side of the characteristic equation (19) for the uncoupled longitudinal mode. Indeed, by using the explicit expressions in Eq.s (13) and (14) for the in-plane and out-of-plane dielectric functions of plasma modes and performing the analytic continuation $i\Omega_{m}\to\omega+i0^{+}$, one can rewrite Eq. (23) as $\varepsilon_{t}(\omega,\eta)=\varepsilon_{\infty}(\omega^{2}-\omega_{ab}^{2})(\omega^{2}-\omega_{c}^{2})/\omega^{2}(\omega^{2}-\omega_{l}^{2}(\eta))$, exactly as in Eq. (8) above.
@@
It is worth mentioning that our derivation is not restricted to electron-doped cuprates, in which the peak has already been experimentally reported [34, 33, 35, 36], but it is in principle valid for any single-layer superconductor, like the hole-doped LSCO. We also point out that the results could be extended to bilayer superconductors like YBCO, that display two Josephson plasmons at frequencies $\omega_{c1}$ and $\omega_{c2}$. Indeed, by using the out-of-plane bilayer dielectric function [5] $\varepsilon_{c}=\varepsilon_{\infty}(\omega^{2}-\omega_{c1}^{2})(\omega^{2}-\omega_{c2}^{2})/\omega^{2}(\omega^{2}-\omega_{T}^{2})$, with $\omega_{T}^{2}=\omega_{c1}^{2}d_{2}+\omega_{c2}^{2}d_{1}$ and $d_{1,2}$ the intra- and inter-bilayer spacings, one predicts two absorptive peaks in the conductivity centered at the high-momenta values of the dispersions of the Josephson modes [19]. The high-energy one follows the same trend of the peak in single-layer superconductors, moving with $\eta$ from $\omega_{c1}$ to $\omega_{ab}$. The low-energy one quickly moves from $\omega_{c2}$ to $\omega_{T}$ even for small values of $\eta$, and does not disappear for $\eta=\pi/2$ [4, 5, 6, 7, 8, 9, 11, 13].

### v2

# Optical absorption in tilted geometries as an indirect measurement of longitudinal plasma waves in layered cuprates

@@ Abstract
Electromagnetic waves propagating in a layered superconductor with arbitrary momentum with respect to the main crystallographic directions display an unavoidable mixing between longitudinal and transverse degrees of freedom. Here we show that this basic physical mechanism explains the emergence of a well-defined absorption peak in the in-plane optical conductivity for light propagating at small tilting angles with respect to the stacking direction in layered cuprates. More specifically, we show that this peak, often interpreted as a spurious leakage of the $c$-axis Josephson plasmon, is instead a signature of the true longitudinal plasma mode occurring at larger momenta. By combining a classical approach based on Maxwell’s equations with a full quantum derivation of the plasma modes based on the modelling of the superconducting phase degrees of freedom, we provide an analytical expression for the absorption peak as a function of the tilting angle and light polarization. We suggest that an all-optical measurement in tilted geometry can be used as an alternative way to access plasma-wave dispersion, usually measured by means of large-momenta scattering techniques like resonant inelastic X-ray scattering (RIXS) or electron energy loss spectroscopy (EELS).
@@ I Introduction
In superconductors the breaking of the continuous gauge symmetry below the superconducting (SC) critical temperature is accompanied by the emergence of two collective modes, associated with the amplitude (Higgs) or phase (Goldstone) fluctuations of the complex SC order parameter, whose absolute value at equilibrium defines the spectral gap for single-particle excitations [1]. While the former is a massive excitation, the latter is massless at long wavelength, reflecting the infinity of possible ground states connected by a global change of the order-parameter phase. Nonetheless, the coupling of the SC phase to the electron density is directly affected by long-range Coulomb interactions between charged electrons. This effect moves the phase mode to the plasma energy scale [2], that is usually much larger than the spectral gap. As a consequence, optical signatures at the plasma energy scale, i.e. at the zero of the dielectric function, are usually unaffected by the SC transition. A rather different phenomenology is instead observed in anisotropic layered superconductors, i.e. systems where the pairing mainly occurs within planes stacked along the $c$ direction, and SC order is established below $T_{c}$ thanks to a weak Josephson-like inter-plane interaction. The hallmark of this category is represented by high-temperature cuprates [3], where the marked anisotropy has been experimentally proven by different optical probes, starting from linear optics, which measures two well-separated energy scales for the plasma modes at long-wavelength for electric fields propagating in the CuO2 planes or perpendicular to them. In these systems the incoherent quasiparticle hopping along the stacking direction makes the $c$-axis response badly metallic: in contrast, below $T_{c}$ the opening of a sizeable spectral gap along with the weak inter-layer pair hopping leave a rather sharp SC plasma edge at a frequency $\omega_{c}$ of few THz in the optical reflectivity, that clearly testifies the emergence of a well-defined SC Josephson plasmon. Even though this feature has been experimentally observed already in the late 1990s [4, 5, 6, 7, 8, 9], renewed interest in the physics of Josephson plasmons emerged more recently. Such interest has been triggered both by the applications to nano-plasmonic [10, 11] and by the role played by Josephson plasmons in non-linear THz spectroscopy [12, 13, 14, 15, 16, 17]. In both cases it becomes theoretically relevant understanding the momentum dependence of the plasmon dispersion at generic momentum, i.e. not along the main crystallographic axes. In this configuration one immediately realizes that the anisotropy leads to a non-trivial response of the system, due to the fact that the current induced by the external electric field is no more parallel to the field itself. As it has been extensively discussed in details in Refs. [18, 19, 20], this mechanism leads to a mixing of the longitudinal and transverse response inside the material, making the distinction between plasmons and polaritons blurred at momenta smaller than a scale $\bar{k}\sim\sqrt{\omega_{ab}^{2}-\omega_{c}^{2}}/c$ set by the anisotropy between in-plane $\omega_{ab}$ and out-of-plane $\omega_{c}$ plasma frequencies. Since usually $\omega_{ab}\gg\omega_{c}$, the effect is relevant for non-linear Josephson plasmonics in the THz regime [21, 22, 23, 24], but does not affect e.g. the measurements of plasmons in RIXS [25, 26, 27, 28, 29] or EELS [30, 31, 32], that usually measure momenta in a fraction of the Brillouin-zone. In the present manuscript we investigate an additional consequence of the above-mentioned mixing, showing how even linear optics can be used to disentangle the longitudinal-transverse mixing in a reflection or transmission geometry which highlights the emergence inside the material of a longitudinal response induced by an external transverse electromagnetic wave. The effect manifests as an absorption peak at a scale nearby $\omega_{c}$ for an electromagnetic wave travelling at small angle with respect to the $c$ direction. This feature has been measured in the past in different samples of electron-doped cuprates [33, 34, 35, 36] below $T_{c}$, and it has been often interpreted as a leakage of the $c$-axis plasmon into the in-plane response [37]. Even more interestingly, the peak position has been shown to change by varying the wave polarization in the plane of incidence, challenging considerably the interpretation of the results. Here we provide a full theoretical description of the microscopic mechanism behind the anomalous absorption peak, and we show that it is a direct consequence of the plasmon-polariton mixing in an anisotropic layered superconductor. We argue that this effect can be used to indirectly probe the plasmon dispersion that usually appears in RIXS and EELS experiments at much larger momenta and, by changing the light polarization, to extract the in-plane and out-of-plane plasma frequencies. Our findings are benchmarked against existing experimental data for cuprates. On a more general ground, our results offer a novel perspective on the possibility to access collective polariton modes in complex materials by properly engineering optical measurements.
@@ II Anisotropic linear response of layered system
$\sigma_{t}(\omega,\eta)=\frac{-\frac{i\omega\varepsilon_{\infty}}{4\pi}\big{(}\sigma_{ab}\cos^{2}{\eta}+\sigma_{c}\sin^{2}\eta\big{)}+\sigma_{ab}\sigma_{c}}{-\frac{i\omega\varepsilon_{\infty}}{4\pi}+\sigma_{ab}\sin^{2}\eta+\sigma_{c}\cos^{2}\eta}.$ (4)
For an electric field polarized along $t$ ($\text{E}_{b}=0$), Eq. (4) immediately gives the conductivity we were looking for, $\sigma_{t}=\text{J}_{t}/\text{E}_{t}$. This expression was first derived in Ref. [34], and its real part displays a peak with central frequency that moves with $\eta$. To show it explicitly, we replace $\sigma_{ab}=\frac{i\omega}{4\pi}\frac{\varepsilon_{\infty}\omega_{ab}^{2}}{(\omega+i0^{+})^{2}}$ and $\sigma_{c}=\frac{i\omega}{4\pi}\frac{\varepsilon_{\infty}\omega_{c}^{2}}{(\omega+i0^{+})^{2}}$, where $\omega_{ab}$ and $\omega_{c}$ are the in-plane and out-of-plane plasma frequencies respectively: one then immediately sees that the real part of $\sigma_{t}(\omega,\eta)$ peaks at frequency $\omega_{l}(\eta)$, that reads
@@
$\varepsilon_{ab}(\omega)=\varepsilon_{\infty}\left(1-\frac{\omega_{ab}^{2}}{(\omega+i0^{+})^{2}}\right),$ (6)
@@
$\varepsilon_{c}(\omega)=\varepsilon_{\infty}\left(1-\frac{\omega_{c}^{2}}{(\omega+i0^{+})^{2}}\right).$ (7)
@@
$\varepsilon_{t}(\omega,\eta)=\varepsilon_{\infty}\frac{\big{(}(\omega+i0^{+})^{2}-\omega_{ab}^{2}\big{)}\big{(}(\omega+i0^{+})^{2}-\omega_{c}^{2}\big{)}}{(\omega+i0^{+})^{2}\big{(}\omega^{2}-\omega^{2}_{l}(\eta)\big{)}}.$ (8)
@@ III.1 Effective action description of plasma modes
To gain a better physical insight into the results of the previous section, we will take advantage of the description of the plasma modes in the SC state obtained via the phase degrees of freedom. Indeed, as it has been recently discussed in Refs. [18, 19], this is a rather powerful and elegant approach to describe the interplay between longitudinal and transverse plasma waves in a layered superconductor, that leads to generalized plasma modes with mixed character at low momenta. We summarize here the main ingredients of the derivation, referring the reader to Refs. [18, 19] and references therein for a detailed derivation of the layered phase-only model.
@@
These gauge-invariant fields provide a full description of the plasma modes once the phase fluctuations are integrated out [18, 19]. To provide simple analytical expressions, in the following, we consider the limit for infinite compressibility. This is a good approximation in single-layer cuprates, as the effects of finite compressibility on the properties of the generalized plasma modes are negligible at small momenta [18]. Interestingly, these effects are actually crucial when studying the optical absorptive peak of bilayer superconductors [4, 5, 6, 7, 8, 9, 13, 15], whose central frequency is significantly influenced by the compressibility [45, 46] due to a capacitive coupling between planes surviving at vanishing momentum [19]. In the basis $\boldsymbol{\psi}_{abc}=\begin{pmatrix}\psi_{a}&\psi_{b}&\psi_{c}\end{pmatrix}^{T}$, and by taking the limit for infinite compressibility that is appropriate for cuprates,
@@
where the plasma frequencies are linked to the in-plane and out-of-plane superfluid stiffness, $\omega_{ab}^{2}=4\pi e^{2}D_{ab}/\varepsilon_{\infty}$ and $\omega_{c}^{2}=4\pi e^{2}D_{c}/\varepsilon_{\infty}$ respectively. Indeed, these go back to Eqs. (6) and (7) once the analytic continuation $i\Omega_{m}\to\omega+i0^{+}$ is performed. Notice that the dielectric tensor is diagonal in the basis $\boldsymbol{\psi}_{abc}$, as $(a,b,c)$ is the reference frame of the principal axes of the crystal. By their definition in Eq. (11), the gauge invariant fields are formally proportional to currents, and we can then apply within the effective-action framework the same procedure used above for the classical approach, i.e. a change of the reference frame to describe a transverse dielectric tensor. We thus perform a rotation around the $b$-axis that combines the $\psi_{a}$ and $\psi_{c}$ components into transverse $\psi_{t}$ and longitudinal $\psi_{l}$ components with respect to the momentum k. The matrix that performs the change of basis $\boldsymbol{\psi}_{abc}\to\boldsymbol{\psi}_{tbl}=\begin{pmatrix}\psi_{t}&\psi_{b}&\psi_{l}\end{pmatrix}^{T}$ reads
@@
where we have performed the analytic continuation $i\Omega_{m}\to\omega+i0^{+}$ and used $k_{c}=|\textbf{k}|\cos\eta$ and $k_{a}=|\textbf{k}|\sin\eta$. Using the definitions of the dielectric functions in Eqs. (13) and (14), the solution of Eq. (19) is exactly the frequency
@@ III.2 Interpretation of the conductivity peak of plasmons
is a dielectric function that describes the transverse linear response of the superconductor along the $t$-axis. Indeed, by making use of the relation $\varepsilon_{\alpha}=\varepsilon_{\infty}+4\pi i\sigma_{\alpha}/\omega$ between the optical conductivity and the dielectric function along the direction $\alpha$ [51] one recovers $\sigma_{t}$ as in Eq. (4). Remarkably, the denominator of $\varepsilon_{t}$ can be brought back to the left-hand side of the characteristic equation (19) for the uncoupled longitudinal mode. Indeed, by using the explicit expressions in Eqs. (13) and (14) for the in-plane and out-of-plane dielectric functions of plasma modes and performing the analytic continuation $i\Omega_{m}\to\omega+i0^{+}$, one can rewrite Eq. (23) as $\varepsilon_{t}(\omega,\eta)=\varepsilon_{\infty}(\omega^{2}-\omega_{ab}^{2})(\omega^{2}-\omega_{c}^{2})/[\omega^{2}(\omega^{2}-\omega_{l}^{2}(\eta))]$, exactly as in Eq. (8) above.
@@
It is worth mentioning that our derivation is not restricted to electron-doped cuprates, in which the peak has already been experimentally reported [34, 33, 35, 36], but it is in principle valid for any single-layer superconductor, like the hole-doped LSCO. We also point out that the results could be extended to bilayer superconductors like YBCO, that display two Josephson plasmons at frequencies $\omega_{c1}$ and $\omega_{c2}$. Indeed, by using the out-of-plane bilayer dielectric function [5] $\varepsilon_{c}=\varepsilon_{\infty}(\omega^{2}-\omega_{c1}^{2})(\omega^{2}-\omega_{c2}^{2})/[\omega^{2}(\omega^{2}-\omega_{T}^{2})]$, with $\omega_{T}^{2}=\omega_{c1}^{2}d_{2}+\omega_{c2}^{2}d_{1}$ and $d_{1,2}$ the intra- and inter-bilayer spacings, one predicts two absorptive peaks in the conductivity centered at the high-momenta values of the dispersions of the Josephson modes [19]. The high-energy one follows the same trend of the peak in single-layer superconductors, moving with $\eta$ from $\omega_{c1}$ to $\omega_{ab}$. The low-energy one quickly moves from $\omega_{c2}$ to $\omega_{T}$ even for small values of $\eta$, and does not disappear for $\eta=\pi/2$ [4, 5, 6, 7, 8, 9, 13, 15]. As mentioned above, finite compressibility corrections are crucial for optical measurements in bilayer superconductors [45, 46], and they must be taken into account to correctly fit the experimental data [19].